# Video Watermarking: Safeguarding Your Video from (Unauthorized) Annotations by Video-based LLMs

### v1

@@ 2.2 Adversarial Attack
While adversarial attacks (Wang et al., 2022; Wei et al., 2019; Bai et al., 2022; Wu et al., 2023; Zou et al., 2023; Zhao et al., 2024; Li et al., 2024a; Gao et al., 2024b, a, c; Fang et al., 2024a, b; Guo et al., 2024; Li et al., 2023a; Xiao et al., 2024; Li et al., 2023c, 2024b) have been widely studied in the context of classification models, the focus of our research lies in the proactive enhancement of video-based LLMs through watermarking. Inspired by the potential vulnerabilities observed in vision tasks, we shift the paradigm to explore how watermarking can fortify large multi-modal models (Qi et al., 2023; Zhuang et al., 2023; Gong et al., 2023; Yang et al., 2024; Zong et al., 2024; Bai et al., 2024; Liang et al., 2024) such as vision large language models (VLLMs) (Zhu et al., 2023; Wang et al., 2024; Liu et al., 2023; Radford et al., 2021b) and text-to-image diffusion models (Rombach et al., 2022). Unlike adversarial attacks designed to manipulate models into generating specific or harmful outputs, our watermarking approach is crafted to protect the integrity and security of video content within these models. In this paper, we introduce a pioneering watermarking strategy that is specifically tailored to safeguard video-based LLMs, ensuring their robustness against unauthorized use and potential data misuse.

### v2

@@ 2.2 Adversarial Attack
While adversarial attacks (Wang et al., 2022; Wei et al., 2019; Bai et al., 2022; Wu et al., 2023; Zou et al., 2023; Zhao et al., 2024; Li et al., 2024a; Gao et al., 2024b, a, c; Fang et al., 2024a, b; Guo et al., 2024; Li et al., 2023a; Xiao et al., 2024; Li et al., 2023c, 2024b; Bai et al., 2019, 2024b, 2020, ; Yang et al., 2024) have been widely studied in classification models, the focus of our research lies in the proactive enhancement of video-based LLMs through watermarking. Inspired by the potential vulnerabilities observed in vision tasks, we shift the paradigm to explore how watermarking can fortify large multi-modal models (Qi et al., 2023; Zhuang et al., 2023; Gong et al., 2023; Yang et al., 2024; Zong et al., 2024; Bai et al., 2024a; Liang et al., 2024) such as vision large language models (VLLMs) (Zhu et al., 2023; Wang et al., 2024; Liu et al., 2023; Radford et al., 2021b) and text-to-image diffusion models (Rombach et al., 2022). Unlike adversarial attacks designed to manipulate models into generating specific outputs, our watermarking approach is crafted to protect the integrity and security of video content within these models. In this paper, we introduce a pioneering watermarking strategy that is specifically tailored to safeguard video-based LLMs, ensuring their robustness against unauthorized use and potential data misuse.